OpenStudy (anonymous):

Katie invested $1200 at 8.5% interest. She earned $612 interest at the end of y years. What is the value of y? A. 18 B. 12 C. 6 D. 3

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

i think it is 12

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@AlexandervonHumboldt2 plz help

OpenStudy (alexandervonhumboldt2):

(si)=ptr/100 si = simple interest 612=si p=1200 r=8.5 (si)=ptr/100 612=1200*t*8.5/100 t=((612*100)/(1200*8.5) now solve the last equation and get the answer.